# Fast and reliable uncertainty quantification with neural network ensembles for industrial image classification

### v2

@@ 3.3 Uncertainty decomposition
Total uncertainty is computed by averaging the predictive distributions (i.e., softmax distributions) over the different members and calculating the entropy $\mathbb{H}$. Aleatoric uncertainty is computed by calculating the entropy in each member and averaging the entropies. Intuitively, aleatoric uncertainty measures uncertainty in the softmax output of individual members; epistemic uncertainty measures how much the members deviate. Naturally, the single NN has zero epistemic uncertainty because it only consists of one member. For a classification problem with $K$ classes, the maximum uncertainty in a prediction is $\log(K)$.
@@ 5.1 Data
The images of the 17 industrial parts are divided into six use cases. The images were generated by rendering 3D CAD models from varying camera angles and under diverse lighting conditions.
@@ 5.4 Evaluation metrics
In addition to the classification accuracy $\uparrow$ (the arrow indicates which direction is better), we examine the predictive uncertainty behavior. Evaluating the predictive uncertainties is challenging as the “ground truth” uncertainty estimates are unavailable. In this work, we use calculate the calibration metric negative log-likelihood (NLL) on the ID set. However, this cannot be calculated on the OOD set because there is no ground truth label and the predictions are incorrect by definition. Uncertainty quantification can be evaluated using classification with rejection and the novel Diversity Quality metric.
@@
The $DQ_{1}$-score is beneficial for practitioners because it simplifies the evaluation of diversity performance. This point is important as the underlying IDD and OODD metrics are difficult to assess separately because they often exhibit a trade-off. Additionally, the $DQ_{1}$-score allows practitioners to consistently prioritize either IDD or OODD by adjusting the $\beta$ value. For instance, in medical applications where incorrect predictions can have severe consequences, practitioners may prioritize achieving high OODD to minimize the risk of incorrect treatments in unseen situations.
@@ 6 Results and Discussion
This section examines the results for both the ID and OOD sets, focusing on predictive performance, uncertainty behavior, Diversity Quality, and computational cost. Finally, we address the trade-offs among these aspects to identify the most suitable NN ensemble for practical applications. We work with a snapshot ensemble of $M=8$ members and a MIMO ensemble of $M=3$ members because this yields the lowest validation NLL for their ensemble type. Please refer to Appendix B for an overview of the validation performance.

### v3

@@ 3.1 Aleatoric and epistemic uncertainty
This study focuses on quantifying aleatoric and epistemic uncertainty in convolutional NNs (CNNs) for image classification. However, these uncertainties can also be computed in other NN architectures, e.g., Transformer-based NNs for text (Thuy et al., 2024), as well as in other machine learning models using alternative techniques, e.g., Gaussian Processes and Random Forests (Hüllermeier and Waegeman, 2021).
@@ 3.3 Uncertainty decomposition
Total uncertainty is computed by averaging the predictive distributions (i.e., softmax distributions) over the different members and calculating the entropy $\mathbb{H}$. Aleatoric uncertainty is computed by calculating the entropy in each member and averaging the entropies. Intuitively, aleatoric uncertainty measures uncertainty in the softmax output of individual members; epistemic uncertainty measures how much the members deviate. Naturally, the single NN has zero epistemic uncertainty because it only consists of one member. For a classification problem with $K$ classes, the maximum total uncertainty in a prediction is $\log(K)$.
@@ 5.1 Data
The dataset authors organized the images into six separate sub-datasets, each representing a specific use case. The first four use cases involve isolated parts requiring classification. Use case 1 includes objects mounted in the cabin of a truck, while use cases 2 to 4 feature parts found in various logistics picking stations. Accurate classification in these cases ensures that the correct parts are either mounted in the truck cabin or selected for delivery. The final two use cases involve assembled objects, where pairs of parts are joined together. The images were generated by rendering 3D CAD models from varying camera angles and under diverse lighting conditions.
@@ 5.4 Evaluation metrics
In addition to the classification accuracy $\uparrow$ (the arrow indicates which direction is better), we examine the predictive uncertainty behavior. Evaluating the predictive uncertainties is challenging as the “ground truth” uncertainty estimates are unavailable. In this work, we calculate the calibration metric negative log-likelihood (NLL) on the ID set. However, this cannot be calculated on the OOD set because there is no ground truth label and the predictions are incorrect by definition. Uncertainty quantification can be evaluated using classification with rejection and the novel Diversity Quality metric.
@@
The $DQ_{1}$-score is a valuable tool for practitioners. It offers a comprehensive evaluation of model diversity by integrating two closely related metrics. While practitioners can assess the IDD and OODD metrics individually, this process is mentally demanding and prone to error.
Moreover, the $DQ_{\beta}$-score is flexible as the $\beta$ parameter can be adjusted based on the specific application, enabling prioritization of one metric over the other. Management can determine an appropriate $\beta$ value by trading off the cost of an incorrect OOD prediction and the cost of consulting a human expert. For instance, in medical applications where incorrect predictions have severe consequences, practitioners can increase the $\beta$ value to emphasize strong OODD performance. Effectively flagging OOD observations with high diversity is essential in such cases, as unseen medical situations demand human intervention for reliable diagnoses.
Additionally, when practitioners assess IDD and OODD metrics separately to guide model selection, they implicitly assume a weighting between these metrics. The $DQ_{\beta}$-score formalizes this weighting through the explicit $\beta$ parameter, resulting in a more transparent and reproducible decision-making process.
@@ 6 Results and Discussion
This section examines the results for both the ID and OOD sets, focusing on predictive performance, uncertainty behavior, Diversity Quality, and computational cost. Finally, we address the trade-offs among these aspects to identify the most suitable NN ensemble for practical applications. We work with a snapshot ensemble of $M=8$ members and a MIMO ensemble of $M=3$ members because this yields the lowest validation NLL for their ensemble type. Please refer to Appendix B for an overview of the validation performance. Furthermore, we set $\beta=1$ in the $DQ_{\beta}$-score and provide an analysis for a varying $\beta$ in Appendix D.